OpenStudy (anonymous):

36 is what % of 450?

OpenStudy (anonymous):

I just posted this question but It isnt coming our right, U r supposed to divide 450 by 36 and it still isnt coming out right for me.

OpenStudy (callisto):

Hmm... 450 (x%) =36 x% = 36/450 x = 36/450 x 100 Can you work it out?

OpenStudy (anonymous):

by what callisto said i got 1250

OpenStudy (callisto):

@suzy4321 I don't agree with you. By what I've said, x would not be greater than 450

OpenStudy (anonymous):

8?

OpenStudy (callisto):

@suzy4321 Yes. Actually, you don't have to leave the answer here. I hope the asker to try, so I didn't give the final answer

OpenStudy (anonymous):

o sorry i was just participating

OpenStudy (kaederfds):

36 = 8 % of 450